A wire has resistance of 88 omega .If it`s length is made half by folding ,find its resistance after folding

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 
Resistance R = 88 ohm
let's take length as L
and cross section area as A

here we need to use the formula R= rho(L/A) 
here rho (as i didn't find the Greek symbol for it) stands for resistivity

as length is halved and r the radius is increased by two 
R is reduced to R/4=88/4=22 ohm

so new resistance will be  22 ohm
 
Regards
Arun (askIITians forum expert)